# Identification of soft modes in amorphous Al2O3 via first-principles

### v1

@@ Abstract
Amorphous Al2O3 is a fundamental component of modern superconducting qubits. While amphorphous oxides offer distinct advantages, such as directional isotropy and a consistent bulk electronic gap, in realistic systems these compounds support two-level systems (TLSs) which couple to the qubit, expediting decoherence. In this work, we perform a first-principles study of amorphous Al2O3 and identify low-energy modes in the electronic and phonon spectra as a possible origin for TLSs.
@@ I Introduction
In recent years, superconducting qubits have emerged as a leading platform for quantum computation in both academia and industry Kjaergaard et al. (2020); Devoret et al. (2004); Koch et al. (2007); Krantz et al. (2019). The existence of two-level systems (TLS) within superconducting qubits have been proposed as a contributing factor to the rate of decoherence Gordon et al. (2014); Bilmes et al. (2017); Müller et al. (2009); Ku and Yu (2005); James et al. (2001); Müller et al. (2019); Mansikkamäki et al. (2024); Crowley et al. (2023); McRae et al. (2020); Gunnarsson et al. (2013); Pritchard and Rondinelli (2024). Despite the presence of TLSs, amorphous Al2O3 remains well suited for use in superconducting qubits, as the lack of long range order has the consequence of making the electronic properties directionally isotropic. Amorphous Al2O3 can be customized to a variety of architectures without concern for grain-boundaries and other complications of cleaving crystalline compoundsMorrissey and Carter (1984), and is dynamically stable with robust electronic properties Lee et al. (2009); Katiyar et al. (2005).
@@
The hypothesis we investigate is that localized low-frequency phonons in amorphous Al2O3 may be considered as TLSs. These “soft localized modes” oscillate with frequencies that are much lower than the typical Bose peak and Debye energies. Hence, once quantized, these modes can induce local TLS dynamics and induce decoherence at low frequencies. This hypothesis has been put forth in other studies however, to the best knowledge of the authors, an attempt to directly compute the phonon modes of amorphous Al2O3 via first-principles has yet to be put forth. To investigate this proposal, we generate multiple amorphous samplesCooper et al. (2000); McCulloch et al. (2000); Štich et al. (1991); Car and Parrinello (1988). The starting point for generation of these samples is supercells of the ordered solid. The ordered solids are then melted using Car-Parinello molecular dynamics Car and Parrinello (1985). Multiple annealing cycles are then carried out until the samples exhibit stability at room temperature. Prior to computation of electronic properties and phonon modes, relaxation of the atomic positions at zero temperature is performed via density functional theory.
@@
The low-frequency phonon modes of the resulting amorphous oxides are computed using density functional perturbation theory and mapped in real-space. Our results provide evidence that, when relaxed at zero-temperature, amorphous Al2O3 support phonon modes at or below 1THz localized in real-space, in alignment with the expectation for a TLS. We chose to examine the window of 0-1 THz rather than the more restricted range of 1-10 GHz at which a qubit generally operates as our first-principles computations encounter finite size effects which can harden soft modes in amorphous solids, shifting frequencies higherWesselinowa and Kovachev (2007). This is discussed as well as potential mitigation strategies through the use of alternate amorphous oxides.
@@ II Generation of amorphous oxides
In order to generate the structures we begin with ordered crystals of Al2O3 in spacegroup R3. A supercell of 120 atoms is formed. These ordered crystals are shown in Fig. (3).
@@ III Low-frequency phonon modes
We now examine the phonon modes of each sample to search for evidence to support or reject the hypothesis that low-frequency modes exist, localized in real space, which serve as TLSs coupling to the qubit. To compute the low-frequency phonon modes we utilize density functional perturbation theory (DFPT) within Quantum Espresso. We limit our computation of phonon modes to the location as the system is not an ordered solid. The lowest-energy phonon modes for each of the samples along with the phonon density of states upon averaging over each sample are shown in Fig. (5).
@@
Examining the density of states we note that the observed, localized low-frequency modes are located primarily in the range THz. It is important to address that this is not the expected frequency of GHz expected for a TLS which couples strongly to the qubit. However, this is not entirely surprising and is likely due in part to finite size effects. Low-frequency modes are long wavelength and may be artificially hardened by size constraints and artifical periodicity invoked by density functional theory. Future work is required to accommodate atoms, creating a unit cell sufficiently large to limit the effects of artificial periodicity. Machine learned interatomic potentials appear ideally positioned to tackle this future challengeChen and Ong (2022); Deng et al. (2023); Batatia et al. (2022).
@@
As TLSs interact with qubits primarily through their electric dipole momentLisenfeld et al. (2015), the dipole moment of each mode is computed and displayed along with the respective phonon mode in Fig. (5). The dipole moment is computed as,
@@
where is the component of the Born effective charge tensor for atom computed via DFT and is similarly the displacement along for atom for a given phonon mode. The expected dipole moment of a TLS in superconducting qubits has been estimated to be Hung et al. (2022). This is in excellent agreement with the dipole moment computed for each of the low-frequency modes, providing evidence towards their classification as dipole activated TLSs.
@@
In order to qualitatively understand the origin of these low-frequency modes we can zoom in on the structure in the region for which the displacement vectors are maximal. Doing so, we observe that these large displacement vectors occur for atoms which exhibit spatial isolation to a greater extent than their neighbors. Due to this spatial isolation we can be consider these states to be akin to “dangling bonds”. An example of this is shown in Fig. (6). This is a unique property of the amorphous systems as an ordered crystalline material will ensure that no atom is isolated in this manner. This observation further provides insight for mitigation of TLSs through increased order.
@@ IV Mitigation strategies
Identification of low-frequency phonon modes as a source of TLSs in superconducting qubits provides an opportunity to determine mitigation strategies. Namely, we have established that these low-frequency phonon modes arise due to the amorphous nature of the system, allowing for the existence of dangling bond-like structures within the bulk of the sample. One route to remove these defects is to work with a compound admitting increased order, such as a quasi-crystal, rather than an amorphous solid. With increased order isolated defects can be minimized.
@@ V Discussion
In this work we have performed detailed first-principles computations to construct multiple samples of amorphous Al2O3. Computing the phonon modes we identify low-frequency states exhibiting real-space localization and a dipole moment in agreement with that expected for TLSs in superconducting qubits which use amorphous Al2O3 as the Josephson junction barrier. We find that the low-frequency modes are inherent in amorphous Al2O3 due to the lack of order and can be mitigated by use of crystalline or quasi-crystalline materials for which such bulk defects are suppressed. The computationally available energies we probe are in sub THz range. We do believe that similar soft localized excitations persist down to GHz range, relevant for device applications, and thus are viable candidates for TLS dephasers in SC qubits. This observation also applies to the existence of TLSs on the surfaces.
@@ Acknowledgements.
- Pritchard and Rondinelli (2024) Pritchard, P. G. and Rondinelli, J. M., “Suppressed paramagnetism in amorphous ta o {} oxides and its link to superconducting qubit performance,” arXiv:2410.13160 (2024), https://doi.org/10.48550/arXiv.2410.13160.
@@ Appendix A Analysis of electronic density of states
In the main body we focus primarily on the identification of soft modes in the phonon spectra as potential two-level systems (TLSs), hastening decoherence in modern superconducting qubits. Computation of the phonon modes was done for bulk samples of amorphous Al2O3 (amorphous Al2O3). This is necessary due to the immense computational expense associated with the generation of large amorphous Al2O3 samples within first-principles, and subsequent computation of the phonon modes. Despite computational considerations limiting analysis of phonon soft modes to the bulk, it is important to consider the surfaces of amorphous Al2O3. Surfaces can support a high-density of dangling bonds manifesting as surface roughness.
Following the observations of dangling bonds in the bulk explored in the main body, dangling bonds on the surface of amorphous Al2O3 are likely candidates for soft phonon modes, i.e. TLSs. Additionally, dangling bonds on the surface can support low-energy electronic bound states. As such, probing the electronic density of states on the surface can be used to determine surface roughness and the density of TLSs. In order to model the surface of amorphous Al2O3 and compute the density of states, we construct real-space tight-binding models for each of the amorphous Al2O3 samples generated in the main body using the Wannier90 software packagePizzi et al. (2020). The bulk electronic states of the tight-binding model precisely matches that computed from first-principles in each case and is shown for the five amorphous samples in Fig. (9). In each case we note the presence of a clean, insulating band-gap in line with expectations from experimental studies.

### v2

@@ Abstract
Amorphous Al2O3 is a fundamental component of modern superconducting qubits. While amorphous oxides offer distinct advantages, such as directional isotropy and a consistent bulk electronic gap, in realistic systems these compounds support two-level systems (TLSs) which couple to the qubit, expediting decoherence. In this work, we perform a first-principles study of amorphous Al2O3 and identify low-energy modes in the electronic and vibrational spectra as a possible origin for TLSs.
@@ I Introduction
In recent years, superconducting qubits have emerged as a leading platform for quantum computation in both academia and industry Kjaergaard et al. (2020); Devoret et al. (2004); Koch et al. (2007); Krantz et al. (2019). The existence of two-level systems (TLS) within superconducting qubits have been proposed as a contributing factor to the rate of decoherence Gordon et al. (2014); Bilmes et al. (2017); Müller et al. (2009); Ku and Yu (2005); James et al. (2001); Müller et al. (2019); Mansikkamäki et al. (2024); Crowley et al. (2023); McRae et al. (2020); Gunnarsson et al. (2013); Pritchard and Rondinelli (2024); Agarwal et al. (2013). Despite the presence of TLSs, amorphous Al2O3 remains well suited for use in superconducting qubits, as the lack of long range order has the consequence of making the electronic properties directionally isotropic. Amorphous Al2O3 can be customized to a variety of architectures without concern for grain-boundaries and other complications of cleaving crystalline compoundsMorrissey and Carter (1984), and is dynamically stable with robust electronic properties Lee et al. (2009); Katiyar et al. (2005).
@@
The hypothesis we investigate is that localized low-frequency phonons in amorphous Al2O3, which we will refer to as vibrons and show schematically in Fig. (1), may be considered as TLSs. These “soft localized modes” oscillate with frequencies that are much lower than the typical Bose peak and Debye energies. Hence, once quantized, these modes can induce local TLS dynamics and induce decoherence at low frequencies. The hypothesis that spatially localized modes, often described as dangling bonds, can give rise to TLSs has been put forth in other studiesBuchenau et al. (1991); Laird and Schober (1991); Wang et al. (2019); Müller et al. (2019) however, to the best knowledge of the authors, an attempt to directly compute the low-frequency modes of amorphous Al2O3 via first-principles has yet to be carried out. To investigate this proposal, we generate multiple amorphous samplesCooper et al. (2000); McCulloch et al. (2000); Štich et al. (1991); Car and Parrinello (1988). The starting point for generation of these samples is supercells of the ordered solid. The ordered solids are then melted using Car-Parinello molecular dynamics Car and Parrinello (1985). Multiple annealing cycles are then carried out until the samples exhibit stability at room temperature. Prior to computation of electronic properties and phonon modes, relaxation of the atomic positions at zero temperature is performed via density functional theory.
@@
The low-frequency phonon modes of the resulting amorphous oxides are computed using density functional perturbation theory and mapped in real-space. Our results provide evidence that, when relaxed at zero-temperature, amorphous Al2O3 support phonon modes at or below 1THz localized in real-space, in alignment with the expectation for a TLS. We chose to examine the window of 0-1 THz rather than the more restricted range of 1-10 GHz at which a qubit generally operates as our first-principles computations encounter finite size effects which can harden soft modes in amorphous solids, shifting frequencies higherWesselinowa and Kovachev (2007). Nevertheless, we can utilize alternative metrics, namely the dipole-moment, to determine the validity of our hypothesis that these modes act as TLSs. Before concluding, we also discuss potential mitigation strategies through the use of alternate amorphous oxides.
@@ II Generation of amorphous oxides
In order to generate the structures we begin with ordered crystals of Al2O3 in spacegroup R3. A supercell of 120 atoms is formed.
@@ III Low-frequency phonon modes
We now examine the phonon spectra of each sample to search for evidence to support or reject the hypothesis that vibrons exist, localized in real space, which serve as TLSs coupling to the qubit. To compute the low-frequency phonon modes we utilize density functional perturbation theory (DFPT) within Quantum Espresso. We limit our computation of phonon modes to the location as the system is not an ordered solid. The lowest-energy phonon modes for each of the samples along with the phonon density of states upon averaging over each sample are shown in Fig. (5).
@@
Examining the density of states we note a peak in the range THz. This peak corresponds directly to the vibron modes for which the displacement vectors are shown in (5)(a)-(e). It is important to address that this is not the expected frequency of GHz expected for a TLS which couples strongly to the qubit. However, this is not entirely surprising and is likely due in part to finite size effects. Low-frequency modes are long wavelength and may be artificially hardened by size constraints and artificial periodicity invoked by density functional theory. Future work is required to accommodate atoms, creating a unit cell sufficiently large to limit the effects of artificial periodicity. Machine learned interatomic potentials appear ideally positioned to tackle this future challengeChen and Ong (2022); Deng et al. (2023); Batatia et al. (2022).
@@
As TLSs interact with qubits primarily through their electric dipole momentLisenfeld et al. (2015), we compute the dipole moment of the identified vibron modes. Experimental studies have determined the dipole moment of TLSs in amorphous aluminum oxide, making it a suitable metric to determine the validity of the computational results. The dipole moment of each mode is computed and displayed along with the respective phonon mode in Fig. (5). The dipole moment is computed as,
@@
where is the component of the Born effective charge tensor for atom computed via DFT and is similarly the displacement along for atom for a given vibron. The expected dipole moment of a TLS in superconducting qubits has been estimated to be Hung et al. (2022). This is in excellent agreement with the dipole moment computed for each of the vibrons, providing evidence towards their classification as dipole activated TLSs.
@@
In order to qualitatively understand the structural origin of these low-frequency modes we can zoom in on the region for which the displacement vectors are maximal. Doing so, we observe that these large displacement vectors occur for atoms which exhibit spatial isolation to a greater extent than their neighbors. To quantify the real-space localization of the low-frequency modes in real-space, we compute the inverse participation ratio as a function of frequency. The inverse participation ratio is computed as,
where for phonon mode , is the normalized displacement vector at atom . If a given phonon mode is localized to a single atom then this value approaches , while a fully delocalized phonon mode yields . We compute the IPR for each phonon mode and amorphous sample. Collecting all values of the IPR as a function of the frequency of the phonon mode, , we can compute a density of states for the IPR as a function of frequency. The results are used to color code the phonon density of states in Fig. (5)(f). This color coding reveals the extent to which the low-frequency modes are localized compared to those at high-frequency. Due to this spatial isolation we can be consider these states to be akin to “dangling bonds”. The existence of dangling bonds within the bulk is a unique property of the amorphous systems as an ordered crystalline material will ensure that no atom is isolated in this manner.
We have therefore successfully computed the low-frequency modes for amorphous samples of Al2O3 and found evidence that vibron modes, appearing as dangling bonds support a dipole moment in agreement with that experimentally measured. However, it is important to briefly discuss the manner in which such dangling bonds can give rise to a TLS. This is explored in Ref. Varma (2023), where it is stated that the spatial isolation of a dangling bond enables a rotational degree of freedom, absent in a periodic crystal. As the environment surrounding the dangling bond is not uniform in the amorphous system, this rotational degree of freedom will produce an anharmonic potential well, providing correspondence between the TLS and dangling bond.
@@ IV Mitigation strategies
Identification of low-frequency modes, vibrons, as a source of TLSs in superconducting qubits provides an opportunity to determine mitigation strategies. Namely, we have established that these vibrons arise due to the amorphous nature of the system, allowing for the existence of dangling bond-like structures within the bulk of the sample. One route to remove these defects is to work with a compound admitting increased order, such as a quasi-crystal, rather than an amorphous solid. With increased order isolated defects can be minimized.
@@ V Discussion
In this work we have performed detailed first-principles computations to construct multiple samples of amorphous Al2O3. Computing the phonon modes we identify low-frequency modes exhibiting real-space localization and a dipole moment in agreement with that expected for TLSs in superconducting qubits which use amorphous Al2O3 as the Josephson junction barrier. Importantly, we find that the low-frequency modes, which we refer to as vibrons, take the form of dangling-bonds and are thus a unique feature of the non-crystalline system. This observation is in-line with prior worksWang et al. (2019); Varma (2023), which have shown that a dangling bond may possesses a rotational degree of freedom not found in an ordered solid. This rotational degree of freedom gives rise to an anharmonic potential well, i.e. TLS. As the dangling bonds arise due to a lack of order, they can be mitigated by use of crystalline or quasi-crystalline materials for which such bulk defects are suppressed. The computationally available energies we probe are in sub THz range. We do believe that similar excitations persist down to GHz range, relevant for device applications, and thus are viable candidates for TLS dephasers in SC qubits. This observation also applies to the existence of TLSs on the surfaces.
@@ Acknowledgements.
- Pritchard and Rondinelli (2024) Pritchard, P. G. and Rondinelli, J. M., “Suppressed paramagnetism in amorphous ta2o5-x oxides and its link to superconducting qubit performance,” arXiv:2410.13160 (2024), https://doi.org/10.48550/arXiv.2410.13160.
- Agarwal et al. (2013) Agarwal, K., Martin, I., Lukin, M. D., and Demler, E., “Polaronic model of two-level systems in amorphous solids,” Phys. Rev. B 87, 144201 (2013).
@@
- Buchenau et al. (1991) Buchenau, U., Galperin, Y. M., Gurevich, V. L., and Schober, H. R., “Anharmonic potentials and vibrational localization in glasses,” Phys. Rev. B 43, 5039–5045 (1991).
- Laird and Schober (1991) Laird, B. B. and Schober, H. R., “Localized low-frequency vibrational modes in a simple model glass,” Phys. Rev. Lett. 66, 636–639 (1991).
- Wang et al. (2019) Wang, L., Ninarello, A., Guan, P., Berthier, L., Szamel, G., and Flenner, E., “Low-frequency vibrational modes of stable glasses,” Nat. commun. 10, 26 (2019).
@@
- Varma (2023) Varma, C. M., “Theory of melting of glasses,” Phil. Mag. 103, 1650–1663 (2023).
@@
- Schofield et al. (2013) Schofield, S., Studer, P., Hirjibehedin, C., Curson, N., Aeppli, G., and Bowler, D., “Quantum engineering at the silicon surface using dangling bonds,” Nat.commun. 4, 1649 (2013).
@@ Appendix A Analysis of electronic density of states
In the main body we focus primarily on the identification of soft modes in the phonon spectra as potential two-level systems (TLSs), hastening decoherence in modern superconducting qubits. Computation of the phonon modes was done for bulk samples of amorphous Al2O3 (amorphous Al2O3). This is necessary due to the immense computational expense associated with the generation of large amorphous Al2O3 samples within first-principles, and subsequent computation of the phonon modes. Despite computational considerations limiting analysis of phonon soft modes to the bulk, it is important to consider the surfaces of amorphous Al2O3. Surfaces can support a high-density of dangling bonds manifesting as surface roughness. The analysis presented in the main body validates the hypothesis that dangling-bonds and similar structurally isolated features in an amorphous system can give rise to localized soft-modes which serve as two-level systems.
Following analysis laid out in the main body, dangling-bind like features are a key structural indicator of localized soft modes which are TLS candidates. Therefore, dangling bonds on the surface of amorphous Al2O3 are likely candidates for soft phonon modes, i.e. TLSs. Additionally, dangling bonds on the surface can support low-energy electronic bound states. This is well studied in prior works on structural disorderSchofield et al. (2013); Wang et al. (2024). As such, probing the electronic density of states on the surface can be used to determine surface roughness and the density of TLSs. In order to model the surface of amorphous Al2O3 and compute the density of states, we construct real-space tight-binding models for each of the amorphous Al2O3 samples generated in the main body using the Wannier90 software packagePizzi et al. (2020). The bulk electronic states of the tight-binding model precisely matches that computed from first-principles in each case and is shown for the five amorphous samples in Fig. (9). In each case we note the presence of a clean, insulating band-gap in line with expectations from experimental studies.